OpenStudy (anonymous):

Find the slope-intercept form for the line satisfying the following conditions. x-intercept 3, y-intercept 4/5 .

OpenStudy (anonymous):



OpenStudy (anonymous):

slope of line is -[4/5]/3 = -4/15 and y-intercept as u know is 4/5 so y = -4x/15 +4/5 is the required form